# Exploring the Interior Structure and Mode of Tidal Heating in Enceladus

### v3

@@ Abstract
Enceladus is among the most intriguing bodies in the solar system due to its astrobiological potential. Determining the extent and duration of habitability (i.e., sustained habitability) at Enceladus requires characterizing its interior properties as well as the level and distribution of tidal heating. Inferring the intensity of geophysical activity in the core has direct implications for understanding the potential for hydrothermal activity and the supply of chemical species important for habitability to the ocean. We build a statistical framework using an MCMC approach to constrain the interior structure based on currently available estimates of libration, shape, heat flux, and total mass. We use this framework to examine the extent to which geodetic measurements can improve our understanding of the interior structure, with an emphasis on determining the partitioning of tidal dissipation between the shell and the core. We quantify plausible ranges of gravitational ($k_{2}$) and displacement ($h_{2}$ and $l_{2}$) tidal Love numbers consistent with existing observations. We demonstrate that measuring the amplitude and phase of $k_{2}$ alone can only constrain the total tidally dissipated energy, but not its radial distribution. However, measuring the amplitude and phase of $h_{2}$ or $l_{2}$ facilitates determining the extent of tidal dissipation in the shell and the core. We provide the precisions required for measuring $k_{2}$, $h_{2}$, and/or $l_{2}$ that enable distinguishing between the two main scenarios of tidal heating, i.e., in the shell versus the core. We also briefly explore the effect of the structural heterogeneities of the shell on the tidal response. Lastly, we evaluate the efficacy of a suite of future geodetic measurements to constrain key interior properties essential to understand the present-day (i.e., instantaneous) and long-term (sustained) habitability at Enceladus.
@@ 1 Introduction and Background
Flybys of Saturn’s moon Enceladus by the Cassini spacecraft revealed multiple intriguing features suggesting an active surface and interior: plumes that continuously erupt from a region approximately centered on the satellite’s south pole (Porco et al., 2006) which feed Saturn’s diffuse E-ring with water-ice crystals (Kempf et al., 2010); a very bright surface (Hendrix et al., 2010); a high surface heat flux (Spencer et al., 2006; Howett et al., 2011); and, importantly, evidence for a subsurface ocean (Porco et al., 2006; Iess et al., 2014; Thomas et al., 2016; Park et al., 2024). This subsurface water body likely meets the criteria for instantaneous habitability (Cockell et al., 2024): An extensive liquid water reservoir containing salts and complex organic compounds and the expected chemical disequilibrium required for supporting metabolism. These conditions can occur through water-rock interactions, presumably in inferred hydrothermal systems at or below the seafloor (Postberg et al., 2009; Howett et al., 2011; Hsu et al., 2015; Cable et al., 2021; Cockell et al., 2024). These lines of evidence for a potentially habitable environment position Enceladus as a compelling target for future exploration by orbiters and landers (NASEM, 2023). Investigating the extent to which these conditions might have been sustained at Enceladus requires understanding the geophysical processes in the interior over geological timescales (see Cockell et al. (2024) and NASEM (2023) for discussions of sustained or dynamical habitability).
@@
Analysis of Cassini’s imaging, thermal mapping, and radiometric tracking data explored the global shape, gravity field, libration (i.e., diurnal wobble of the ice shell), and the surface topography of Enceladus, providing insights into the interior layers’ properties. The large amplitude of forced physical libration suggests that the ice crust is mechanically decoupled from the rocky core, likely by a global ocean (Nadezhdina et al., 2016; Thomas et al., 2016; Van Hoolst et al., 2016; Park et al., 2024). The presence of a global ocean is also supported by analysis of the low-order gravity field (McKinnon, 2015; Čadek et al., 2016).
@@
The total rate of heat loss at Enceladus remains poorly constrained (Howett et al., 2025). Estimates of the ice shell thickness and surface temperatures suggest a total heat loss of approximately 25-40 GW (Hemingway & Mittal, 2019; Nimmo et al., 2023), while the measured heat output from the South Polar Terrain (SPT) alone is estimated to be 4-19 GW (Howett et al., 2011). With an estimated radiogenic heating rate of less than 0.3 GW (Vance et al., 2007; Roberts & Nimmo, 2008; Nimmo et al., 2018) and minimal additional heating due to libration (Shao & Nimmo, 2022), the most likely heat source is therefore dissipation due to the eccentricity tides exerted by Saturn (e.g., Nimmo et al., 2023). Initial analyses suggested that given the global properties of Enceladus and Saturn, tidal dissipation is insufficient to explain the estimated heat loss (Meyer & Wisdom, 2007). However, later studies suggested that Saturn is supplying tidal energy and angular momentum to its satellites at a higher rate than previous estimates (Lainey et al., 2017), suggesting a resonance locking between the moons and the planet’s internal oscillation modes (Fuller et al., 2016; Lainey et al., 2017; Nimmo et al., 2018; Lainey et al., 2020; Jacobson, 2022). Alternatively, there may exist transient periods of tidal heating. In this scenario, the present-day ocean would be a manifestation of a cycle of ocean formation and freezing as a result of a dynamically unstable orbit (Meyer & Wisdom, 2008a; Goldreich et al., 2025), or instead as a result of changes in the mean motion resonance in the case of an ongoing feedback between the moon’s thermal and orbital evolution. (Ojakangas & Stevenson, 1986; Meyer & Wisdom, 2008b; Renaud et al., 2021; Goldreich et al., 2025). A transient ocean could also be the result of recent catastrophic events that affect the Saturnian system involving Enceladus (e.g., Asphaug & Reufer, 2013; Wisdom et al., 2022; Teodoro et al., 2023).
@@
Considering the ambiguities in understanding the interior, geophysical investigations of Enceladus must account for the possibility of heating in the deep interior. Ermakov et al. (2021) and Genova et al. (2024) explore models of the interior by measuring the potential tidal Love number $k_{2}$ along with the libration or gravity-topography admittance to constrain the interior structure (i.e., thicknesses and densities of layers) at Enceladus. However, the ranges of interior parameters considered in these studies assume a nearly rigid core and thus inherently produce negligible tidal dissipation in the core. Therefore, these studies do not address the challenge of discriminating between the different scenarios of tidal dissipation. As we will show, a tidally active core will affect the time-dependent gravity field of the body and produce a gravitational signal that resembles that of a tidally dissipative ice shell, resulting in a non-unique interpretation of the amplitude and phase of $k_{2}$. In addition, gravity and topography measurements have revealed structural heterogeneities in the Enceladus ice shell (Yin & Pappalardo, 2015; Park et al., 2024; Schenk & McKinnon, 2024). Deviations of the body from spherical symmetry can potentially affect its tidal response (e.g., Zhong et al., 2012; Qin et al., 2014; Wahr et al., 2014; Lau et al., 2015; Dmitrovskii et al., 2022; Berne et al., 2023a) by driving deformation at higher harmonic degrees (e.g., $l=3,4$) through mode coupling which is expected to produce variations in degree-2 Love numbers evaluated at different spherical harmonic orders, i.e., $k_{2m}$ and $h_{2m}$ (e.g., Bagheri et al., 2021; Pou et al., 2022; Berne et al., 2023a).
Here, we review the physical aspects of the libration and tidal responses of Enceladus as a generic icy satellite. We explore the range of the interior model parameters compatible with the currently available observations using an MCMC statistical framework. We use this framework to examine the effectiveness of a suite of geodetic measurements including static gravity field, complex tidal Love numbers, libration, and heat loss rate that can constrain the interior of Enceladus.
We highlight the degeneracy in interpreting $k_{2}$ to identify the tidally active region(s) of Enceladus and provide a roadmap to distinguish between the two classes of tidal dissipation. We use our framework to compute the required precisions for measuring the tidal and displacement Love numbers. This framework allows for an improved understanding of the sensitivity of the observables to each of the interior parameters. Using the proposed suite of future geophysical measurements and synthetic observations that could be available in future missions, we consider constraining interior parameters, such as the thickness, density, and rheology of the interior layers. These constraints can significantly improve our understanding of the parameters that control the instantaneous and sustained habitability at Enceladus. The methodology and derivations developed here apply broadly to other icy bodies.
@@ 2 Methods
We describe the geophysical models needed to compute the geodetic quantities in our forward model. The models include the tidal response in the form of the complex Love numbers and maximum surface deformations, the viscoelastic models used in computing the tidal response and physical libration, and calculation of the static gravity coefficients. Detailed derivations are provided in the appendices.
@@ 2.1 Tidal Response
The eccentric orbit of Enceladus around Saturn drives periodic gravitational tidal forcing of the moon. This forcing potential, truncated at spherical harmonic degree 2 and to first order in eccentricity, can be written as (Murray & Dermott, 1999):
@@
where $r$ is the radius, $\theta$ is the co-latitude, $\phi$ is the longitude, $t$ is time, $\omega$ is the frequency of the forcing tide equal to the orbital frequency 53.07 $\mu\rm Rad.s^{-1}$, $e$ is the orbital eccentricity equal to 0.0047, $P_{2}^{0}$ and $P_{2}^{2}$, are the un-normalized associated Legendre polynomials, and $\lambda=\cos\theta$. For a spherically symmetric body, the tidal deformations in spherical coordinates in radial and tangential directions are described by the displacement Love numbers $h_{2}$ and $l_{2}$ (Melchior, 1983):
@@
$k_{2}$, $h_{2}$, and $l_{2}$ are sensitive to the properties of the entire body and primarily depend on the density and rheological structure. The maximum displacements in the radial and horizontal directions are expressed as (Park et al., 2020):
@@
In general, for a body with lateral variations in structure, the spherical harmonic expansion of the tidal response depends on both the degree (here $l=2$) and the order $m$. Thus, whereas for a spherically symmetric body, $k_{22}=k_{20}=k_{2}$ and $h_{22}=h_{20}=h_{2}$, for a non-spherically symmetric body these quantities can be different. In this case, $k_{2m}=W_{2m}(r,\theta,\phi,t)/U_{2m}(r,\theta,\phi,t)$, and $h_{2m}=gH_{2m}(r,\theta,\phi,t)/U_{2m}(r,\theta,\phi,t)$, where $W_{2m}$ and $H_{2m}$ are the spherical harmonic expansion terms of the potential and surface displacement (assuming l=2 forcing only).
@@
Here, $\epsilon_{k_{2}}$ is the phase lag of the perturbation in the gravitational potential, which can represent the total tidal energy generated throughout the body. In most cases, the tidal quality factor $Q$ is simply defined as $\tan(\epsilon_{k_{2}})$. However, this simplification is only correct when the phase lag is small such that $\tan(\epsilon_{k_{2}})\approx\sin(\epsilon_{k_{2}})$. For bodies like Enceladus that may experience intense tidal heating and high tidal phase lag, this approximation is no longer valid (Efroimsky, 2012). Here, we explicitly consider phase lags, as opposed to the quality factor $Q$.
@@ 2.2 Viscoelastic Model
To predict the tidal response of Enceladus, we must consider rheological models that incorporate shear dissipation in rock and icy regions. When both viscous and elastic mechanisms are active, the viscoelastic response depends on the shear modulus $\mu$, frequency of excitation $\omega$, and the effective viscosity $\eta$. The Maxwell viscoelastic model provides a reasonable description of a material’s behavior at time scales that are significantly shorter or longer than the Maxwell (relaxation) time of the body defined as $\tau_{M}=\eta/\mu$. $\tau_{M}$ represents the approximate timescale at which the behavior transitions from being dominated by the elastic regime to the viscous regime. However, a Maxwell model may not be appropriate for describing tidal dissipation in planetary bodies that have tidal periods ranging from hours to a few days, which may be near the relaxation time of the body. Thus, applicability of Maxwell’s model for capturing tidal dissipation has been questioned for multiple planetary bodies (Bills et al., 2005; Castillo-Rogez et al., 2011; Bagheri et al., 2019; Petricca et al., 2024). Several models based on micro-scale physical processes and laboratory measurements have been proposed to mimic the behavior of a material in time scales between the elastic and viscous regimes (e.g., Castillo-Rogez et al., 2011; Bagheri et al., 2019, 2021; Petricca et al., 2024; Tobie et al., 2025). In particular, the Andrade model has been favored to model tidal dissipation in celestial bodies taking into account the effect of anelastic behavior in the transition from dominantly elastic to dominantly viscous regimes (Castillo-Rogez et al., 2011; Renaud & Henning, 2018; Padovan et al., 2014; Bagheri et al., 2019, 2022a). A detailed discussion of various viscoelastic models can be found in Jackson & Faul (2010); Faul & Jackson (2015); Bagheri et al. (2019). We consider both Maxwell and Andrade viscoelastic models because of the potential proximity of the relaxation time of the ice shell to Enceladus’s orbital period, but most of our results assume the Andrade rheology.

### v5

@@ Abstract
Enceladus is among the most intriguing bodies in the solar system due to its astrobiological potential. Determining the extent and duration of habitability (i.e., sustained habitability) requires characterizing the interior properties and the level and distribution of tidal heating in Enceladus. Inferring the intensity of geophysical activity in the core has direct implications for the potential hydrothermal activity and supply of chemical species important for habitability to the ocean. We build a statistical framework to constrain the interior using estimates of libration, shape, heat flux, gravity, and total mass. We use this framework to examine the extent that geodetic measurements can improve our understanding of the interior structure, with an emphasis on partitioning of dissipation between the shell and the core. We quantify plausible ranges of gravitational ($k_{2}$) and displacement ($h_{2}$, $l_{2}$) tidal Love numbers consistent with existing observations. We demonstrate that measuring $k_{2}$ alone can only constrain the total tidally dissipated energy, but not its radial distribution. However, measuring the amplitude and phase of $h_{2}$ or $l_{2}$ facilitates determining the extent of tidal dissipation in the shell and the core. We provide the precisions required for measuring $k_{2}$, $h_{2}$, and $l_{2}$ that enable distinguishing between the main tidal heating scenarios, i.e., in the shell versus the core. We also explore the effect of the structural heterogeneities of the shell on the tidal response. Lastly, we evaluate the efficacy of future geodetic measurements to constrain key interior properties essential to understand the present-day (instantaneous) and long-term (sustained) habitability at Enceladus.
@@ 1 Introduction and Background
Flybys of Saturn’s moon Enceladus by the Cassini spacecraft revealed multiple intriguing features suggesting an active surface and interior: jets that continuously erupt from a region centered on the satellite’s south pole (Porco et al., 2006) and feed Saturn’s diffuse E-ring with water-ice crystals (Kempf et al., 2010); a very bright surface (Hendrix et al., 2010); a high surface heat flux (Spencer et al., 2006; Howett et al., 2011); and, importantly, evidence for a subsurface ocean (Porco et al., 2006; Iess et al., 2014; Thomas et al., 2016; Park et al., 2024). This subsurface water body likely meets the criteria for instantaneous habitability (Cockell et al., 2024): An extensive liquid water reservoir containing salts and complex organic compounds and the expected chemical disequilibrium required for supporting metabolism. These conditions can occur through water-rock interactions, presumably in the inferred hydrothermal systems at or below the seafloor (Postberg et al., 2009; Howett et al., 2011; Hsu et al., 2015; Cable et al., 2021; Cockell et al., 2024). These lines of evidence for a potentially habitable environment position Enceladus as a compelling target for future exploration by orbiters and landers (NASEM, 2023). Investigating the extent to which these conditions might have been sustained at Enceladus requires understanding the geophysical processes in the interior over geological timescales (see Cockell et al. (2024) and NASEM (2023) for discussions of sustained or dynamical habitability).
@@
Analysis of Cassini’s imaging, thermal mapping, and radiometric tracking data explored the global shape, gravity field, libration (i.e., diurnal wobble of the ice shell), and the surface topography of Enceladus, providing insights into the interior properties. The large amplitude of forced physical libration suggests that the ice crust is mechanically decoupled from the rocky core, likely by a global ocean (Nadezhdina et al., 2016; Thomas et al., 2016; Van Hoolst et al., 2016; Park et al., 2024). The presence of a global ocean is also supported by the analysis of the low-order gravity field (McKinnon, 2015; Čadek et al., 2016).
@@
The total rate of heat loss at Enceladus remains poorly constrained (Howett et al., 2025). Estimates of the ice shell thickness and surface temperatures suggest a total heat loss of approximately 25-40 GW (Hemingway & Mittal, 2019; Nimmo et al., 2023), while the measured heat output from the South Polar Terrain (SPT) alone is estimated to be 4-19 GW (Howett et al., 2011). With an estimated radiogenic heating rate of less than 0.3 GW (Vance et al., 2007; Roberts & Nimmo, 2008; Nimmo et al., 2018) and minimal additional heating due to libration (Shao & Nimmo, 2022), the most likely heat source is therefore dissipation due to the eccentricity tides exerted by Saturn (e.g., Nimmo et al., 2023). Initial analyses suggested that given the orbital configuration of Enceladus and Saturn, tidal dissipation is insufficient to explain the estimated heat loss (Meyer & Wisdom, 2007). However, later studies suggested that Saturn is supplying tidal energy and angular momentum to its satellites at a higher rate than previous estimates (Lainey et al., 2017), suggesting a resonance locking between the moons and the planet’s internal oscillation modes (Fuller et al., 2016; Lainey et al., 2017; Nimmo et al., 2018; Lainey et al., 2020; Jacobson, 2022). Alternatively, there may exist transient periods of tidal heating. In this case, the present-day ocean would be a manifestation of a cycle of ocean formation and freezing as a result of a dynamically unstable orbit (Meyer & Wisdom, 2008a; Goldreich et al., 2025), or instead as a result of changes in the mean motion resonance in the case of an ongoing feedback between the moon’s thermal and orbital evolution. (Ojakangas & Stevenson, 1986; Meyer & Wisdom, 2008b; Renaud et al., 2021; Goldreich et al., 2025). A transient ocean could also be the result of recent catastrophic events that affect the Saturnian system (e.g., Asphaug & Reufer, 2013; Wisdom et al., 2022; Teodoro et al., 2023).
@@
Considering the ambiguities in understanding the interior, geophysical investigations of Enceladus must account for the possibility of heating in the deep interior. Ermakov et al. (2021) and Genova et al. (2024) explore models of the interior by future measurements the potential tidal Love number $k_{2}$ along with libration and gravity-topography admittance to constrain the interior structure (i.e., thicknesses and densities of layers) at Enceladus. However, the ranges of interior parameters considered in these studies assume a nearly rigid core and thus inherently produce negligible tidal dissipation in the core. Therefore, these studies do not address the challenge of discriminating between the different scenarios of tidal dissipation. As we will show, a tidally active core will affect the time-dependent gravity field of the body and produce a gravitational signal that resembles that of a tidally dissipative ice shell, resulting in a non-unique interpretation of the amplitude and phase of $k_{2}$. In addition, gravity and topography measurements have revealed structural heterogeneities in the Enceladus ice shell (Yin & Pappalardo, 2015; Park et al., 2024; Schenk & McKinnon, 2024). Structural heterogeneities in the shell can potentially affect its tidal response compared to a spherically symmetric body (e.g., Zhong et al., 2012; Qin et al., 2014; Wahr et al., 2014; Lau et al., 2015; Dmitrovskii et al., 2022; Berne et al., 2023a) by driving deformation at higher harmonic degrees (e.g., $l=3,4$) through mode coupling. Such a coupling can produce variations in degree-2 Love numbers evaluated at different spherical harmonic orders, i.e., $k_{2m}$ and $h_{2m}$ (e.g., Bagheri et al., 2021; Pou et al., 2022; Berne et al., 2023a).
The next sections of the paper are organized as follows. In Section 2, we define the geophysical observables used in this study—tidal response, layer-resolved tidal heating, physical libration, and gravity coefficients—and detail their computation. In Section 3, we analyze the sensitivity of physical libration to shell properties and its correlation with the tidal response. In Section 4, we explore tidal-dissipation regimes in the shell and core that are consistent with the inferred heat-production rate, and we delineate the associated ranges of tidal Love numbers and rheological parameters. In Section 5, we present a Bayesian MCMC framework that integrates all available observations to infer plausible interior models and to estimate ranges for as-yet unmeasured tidal Love numbers. We show the degeneracy of using $k_{2}$ alone to identify where tidal dissipation occurs at Enceladus, assess whether joint observations of $k_{2}$ and $h_{2}$ can distinguish dissipation regimes, and derive the minimum measurement precision required. Using precisions that can be expected from future missions, we perform inversions of geodetic data (static gravity, complex tidal Love numbers, and libration) to evaluate their ability to constrain interior structure. Finally, in Section 6, we summarize the main conclusions and discuss the limitations of our analysis.
@@ 2 Methods
We describe the geophysical models needed to compute the geodetic quantities in our forward calculation. We describe the tidal response in the form of the complex Love numbers and maximum surface deformations and the viscoelastic models used in computing the tidal response, followed by our derivation for the tidal heating in the shell and in the core. We also describe modeling the physical libration, and calculation of the static gravity coefficients. Detailed derivations are provided in the appendices.
@@ 2.1 Tidal Response
The eccentric orbit of Enceladus around Saturn drives periodic gravitational tidal forcing of the moon. The forcing potential, truncated at spherical harmonic degree 2 and to first order in eccentricity, can be written as (Murray & Dermott, 1999):
@@
where $r$ is the radius, $\theta$ is the co-latitude, $\phi$ is the longitude, $t$ is time, $\omega$ is the frequency of the forcing tide equal to the orbital frequency 53.07 $\mu\rm Rad.s^{-1}$, $e$ is the orbital eccentricity equal to 0.0047, $P_{2}^{0}$ and $P_{2}^{2}$ are the un-normalized associated Legendre polynomials, and $\lambda=\cos\theta$. For a spherically symmetric body, the tidal deformations in radial and tangential directions are described by the degree 2 displacement Love numbers $h_{2}$ and $l_{2}$ (Melchior, 1983):
@@
$k_{2}$, $h_{2}$, and $l_{2}$ are sensitive to the global scale properties of the body and primarily depend on its density and rheological structure. The maximum displacements in the radial and horizontal directions are expressed as (Park et al., 2020):
@@
In general, for a body with lateral variations in structure, the spherical harmonic expansion of the tidal response depends on both the degree (here $l=2$) and the order $m$. Thus, whereas for a spherically symmetric body, $k_{22}=k_{20}=k_{2}$ and $h_{22}=h_{20}=h_{2}$, for a non-spherically symmetric body these quantities can be different from each other. In this case, $k_{2m}=W_{2m}(r,\theta,\phi,t)/U_{2m}(r,\theta,\phi,t)$, and $h_{2m}=gH_{2m}(r,\theta,\phi,t)/U_{2m}(r,\theta,\phi,t)$, where $W_{2m}$ and $H_{2m}$ are the spherical harmonic expansion terms of the potential and surface displacement (assuming l=2 forcing only).
@@
Here, $\epsilon_{k_{2}}$ is the phase lag of the perturbation in the gravitational potential. In most cases, the tidal quality factor $Q$ is simply defined as $\tan(\epsilon_{k_{2}})$. However, this simplification is only correct when the phase lag is small such that $\tan(\epsilon_{k_{2}})\approx\sin(\epsilon_{k_{2}})$. For bodies like Enceladus that may experience intense tidal heating and high tidal phase lag, this approximation is no longer valid (Efroimsky, 2012). Here, we explicitly consider phase lags, as opposed to the quality factor $Q$.
@@ 2.2 Viscoelastic Model
To predict the tidal response of Enceladus, we must consider rheological models that quantify shear dissipation in rock and icy regions. When both viscous and elastic mechanisms are active, the viscoelastic response depends on the shear modulus $\mu$, frequency of excitation $\omega$, and the effective viscosity $\eta$. The Maxwell viscoelastic model provides a reasonable description of a material’s behavior at timescales that are significantly shorter or longer than the Maxwell (relaxation) time of the body defined as $\tau_{M}=\eta/\mu$. $\tau_{M}$ represents the approximate timescale at which the behavior transitions from being dominated by the elastic regime to the viscous regime. However, a Maxwell model may not be appropriate for describing tidal dissipation in planetary bodies that have tidal periods ranging from hours to a few days, which may be near the relaxation time of the body. Applicability of Maxwell’s model for capturing tidal dissipation has been questioned for multiple planetary bodies (Bills et al., 2005; Castillo-Rogez et al., 2011; Bagheri et al., 2019; Petricca et al., 2024). Several models based on micro-scale physical processes and laboratory measurements have been proposed to mimic the behavior of a material in time scales between the elastic and viscous regimes (e.g., Castillo-Rogez et al., 2011; Bagheri et al., 2019, 2021; Petricca et al., 2024; Tobie et al., 2025). In particular, the Andrade model has been favored to model tidal dissipation in celestial bodies taking into account the effect of anelastic behavior in the transition from dominantly elastic to dominantly viscous regimes (Castillo-Rogez et al., 2011; Renaud & Henning, 2018; Padovan et al., 2014; Bagheri et al., 2019, 2022a). A detailed discussion of viscoelastic models can be found in Jackson & Faul (2010); Faul & Jackson (2015); Bagheri et al. (2019). We consider both Maxwell and Andrade viscoelastic models because of the potential proximity of the relaxation time of the ice shell to Enceladus’s orbital period (see Appendix A), but most of our results assume the Andrade rheology.
2.3 Partitioning of tidal heating between the shell and the core
The total tidal energy dissipated in the body can be computed by measuring the amplitude and phase lag of $k_{2}$, independent of the choice of rheological model (see equation (5)). We can compute the tidal heating in a specific region of the satellite, such as the core or the shell, by integrating the radial shear dissipation in that part of the body. Thus, if we compute the contribution of each of the layers to $|k_{2}|/Q$ or identically to $\Im(k_{2})$ of the body, that is, evaluating at each interface, we can compute the associated tidal heating for the part of the body that is interior to the interface. Here, we use equations (33), (35), and (37) in Tobie et al. (2005) to derive the contribution of the core and the shell to the global tidal heating:
Here, $\hat{k}_{2}^{core}$ is the piecewise gravitational Love number associated with the core. For a homogeneous core, $\hat{k}_{2}^{core}$ can be computed from the deformation at the top of the core ($\Delta R_{core}$):
$\Im{(\hat{k}_{2}^{core}})=\frac{4\pi G\rho_{core}R_{core}^{4}}{5R_{E}^{3}}\Im{(\Delta R_{core})^{4}}.$ (8)
Details of this derivation are presented in Appendix B. The tidal heating in the ice shell is then calculated by subtracting the contribution of the core from the total energy dissipated:
$\dot{E}_{shell}=\dot{E}-\dot{E}_{core}.$ (9)
Tidal dissipation in the ocean is assumed negligible (Hay & Matsuyama, 2017; Matsuyama et al., 2018). Equations (7) and (9) provide a method to calculate the partitioning of the global tidal heating for a given interior structure model. Thus, these equations require a priori assumptions or independent constraints on the rheological and structural properties to compute the shear dissipation in the core. Here, we derive analytical expressions for the partitioning of tidal heating on the basis of the observable quantities, i.e., the tidal Love numbers and the density structure of the body. We extend the formulation provided by Beuthe (2019) which uses membrane mechanics to obtain the contribution of each layer to the imaginary part of the tidal Love number. The shell’s contribution can be written as (Beuthe, 2019):
where, following the terminology used in Beuthe (2019), $\xi_{2}=3\rho_{ocean}/5\rho_{E}$, $\rho_{E}$ is the density of the whole body, $k_{2}$ and $h_{2}$ are the tidal Love numbers of the whole body, $k_{2}^{0}$ is the fluid-crust gravitational Love number (i.e., assuming the crust is fluid), and $\Lambda_{T}$ is the shell spring constant introduced by Beuthe (2019). In Appendix B, we demonstrate that the tidal dissipation occurring in the shell can be expressed as:
Here, $\delta\rho$ is the density difference between the ocean and the crust and $\epsilon$ is the ratio of the thickness of the crust to Enceladus’s radius. Equation (11) expresses the tidal heating in the shell using measurements of the amplitudes and phases of the Love numbers ($k_{2}$, $h_{2}$, $\epsilon_{k_{2}}$, $\epsilon_{h_{2}}$), and information on the density structure. The densities of the layers can be constrained by the static gravity field and libration, and are easier to constrain compared to the rheology of the layers required in Equation (10). Based on Equation (11), the uncertainty in the magnitude of tidal heating occurring in the shell can be computed from the uncertainties on the density structure and the complex Love numbers. The tidal heating in the core can be computed by subtracting the tidal dissipation in the shell from the total heat. This derivation is general and can be used for other icy satellites for which the necessary geodetic observations are or will become available.